OpenStudy (anonymous):

What is the limiting reactant when 4.5 moles of aluminum react with 6.7 moles of oxygen gas? (4 points) Unbalanced equation: Al + O2 → Al2O3 @TheSmartOne

TheSmartOne (thesmartone):

Can you first find the balanced equation?

TheSmartOne (thesmartone):

Correct :)

TheSmartOne (thesmartone):

Well, so do you see how there are less moles of \(\sf O_2\) that react with \(\sf Al\)? That means that there will be an excess of \(\sf O_2\) because there are more moles available/

OpenStudy (anonymous):

So Al is the limiting reactant?

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

Yes :)

TheSmartOne (thesmartone):

Do you want to know the math way too?

OpenStudy (anonymous):

For educational purposes, yes.

TheSmartOne (thesmartone):

\[\sf \frac{ 4.5 ~mol ~Al}{ 1 } \times \frac{ 3~mol~O_2 }{ 4~mol Al }= 3.2~ mol~O_2\] And since we have been given 6.7 mols but we are only using 3.2 moles, we can obviously say that \(\sf O_2\) is in excess.

TheSmartOne (thesmartone):

And that \(\sf Al\) is the limiting reactant \(\LARGE\ddot\smile\)

OpenStudy (anonymous):

Thanks a lot!!

OpenStudy (anonymous):

@TheSmartOne Where did you get the 4 molAl?

OpenStudy (anonymous):

I need to know this for another question. @TheSmartOne

TheSmartOne (thesmartone):

Sorry was afk.

TheSmartOne (thesmartone):

@brostep0s I took it from the balanced equation